OpenStudy (hollz32):

What is the general form of the equation of a circle with its center at (-2, 1) and passing through (-4, 1)? x2 + y2 − 4x + 2y + 1 = 0 x2 + y2 + 4x − 2y + 1 = 0 x2 + y2 + 4x − 2y + 9 = 0 x2 − y2 + 2x + y + 1 = 0

OpenStudy (welshfella):

you need to find the radius of the circle 0 oh r^2 i mean ( r^2 = (y2 - y1)^2 + (x2-x1)^2 where (x1,y1) and (x2,y2) are the corodinates of the center and the point on the circl

OpenStudy (welshfella):

here it is (-2-(-4))^2 + (1-1)^2

OpenStudy (welshfella):

can you work that out/

OpenStudy (hollz32):

so they answer I get would be the radius? @welshfella

OpenStudy (welshfella):

no it will be the square of the radius which is what we want - its really pythagoras theorum that we are using

OpenStudy (hollz32):

okay. i got 4 for the answer. what would I do to get the answer. like what steps? i'd like to try to understand it.

OpenStudy (hollz32):

would the new equation use the square of the radius? like 4^2=(1-1)^2+(-4-2)^2 @welshfella

OpenStudy (k12awesomeness):

one year ago (x - h)^2 + (y - k)^2 = r^2 ; (h,k) = (-2,1) and r is the distance between (-2,1) and (-4,1) So then plug in the values given, expand the squared expressions and subtract r^2 from both sides of the equation to get your equation in general form.

OpenStudy (k12awesomeness):

Distance between (-2,1) and (-4,1) is 2. (x - (-2))^2 + (y - 1)^2 = (2)^2 --> (x+2)^2 + (y-1)^2 = 4 --> x^2 + 4x + 4 + y^2 - 2y + 1 = 4 --> x^2 + 4x + y^2 - 2y + 5 = 4 --> x^2 + 4x + y^2 - 2y + 1 = 0

OpenStudy (k12awesomeness):

Or you can do x^2 + y^2 + 4x - 2y + 1 = 0

OpenStudy (k12awesomeness):

make sense?

OpenStudy (hollz32):

so I'd subtract the actual; distance from what part of the equation?

OpenStudy (decarr432):

Math is not my best I apologize for not being of any help